Assume that a password to log onto a computer account is to consist of 2 letters and 2 digits, Determine the number of possible in the case that repetition is permitted as well as not permitted.
Answer by sudhanshu_kmr(1152)


As i got that firstly 2 letters and then 2 digits..

number of ways when repetition is permitted = 26*26*10*10 = 67600

number of ways when repetition not allowed = 26*25*10*9 = 58500

if any doubt, feel free to contact..